# Energy-Aware Lane Planning for Connected Electric Vehicles in Urban Traffic: Design and Vehicle-in-the-Loop Validation

### v1

@@ VI Experiments
In this section, we validate our proposed lane selection algorithm on a test vehicle. Further, we compare the energy consumption of the proposed lane selection algorithm with a human driver in the same traffic scenario using the vehicle-in-the-loop testing platform.
@@ VI-1 Experimental Setup
We developed a Vehicle-in-the-Loop (VIL) system that enables safe and efficient testing of autonomous driving algorithms by integrating a Hyundai IONIQ 5 as a physical test vehicle with the CARLA simulator, which generates realistic virtual scenarios including road networks, traffic participants, and traffic lights. The system features a real-time interaction between the actual vehicle and the virtual environment, allowing an accurate replication of traffic conditions and responsive behaviors. Fig. 5 illustrates the virtual CARLA simulator map, the satellite image of the actual testing site, and the physical test vehicle (IONIQ 5). This setup enables the ego vehicle to perceive and react to virtual agents and infrastructure as if operating in the real world, allowing for robust evaluation of planning and control algorithms in a safe environment while maintaining real energy measurements. Additional details on the hardware and simulation setup are provided in [12].
@@
Each experiment consists of three laps around the route (approximately 4 km). To evaluate the proposed algorithms, we designed two baseline approaches and conducted experiments to collect the corresponding energy consumption measurements. The first baseline, referred to as Baseline 1, involves a human driver operating the vehicle while observing the CARLA virtual environment and making lane changes as they would in a real-world setting. The second baseline, referred to as Baseline 2, is an autonomous energy-efficient control strategy that does not perform lane changes, which is essentially a lane-keeping algorithm that optimizes vehicle behavior based on an energy stage cost.
@@ VI-2 Experiment Results
A total of 10 tests were conducted. Fig. 6(a) presents the total motion energy consumption, calculated by subtracting auxiliary power usage from the overall energy consumption, while Fig. 6(b) shows the total energy usage including auxiliary loads. The human driver (Baseline 1) tended to drive closer to the speed limit and changed lanes as needed, resulting in the shortest average trip time of 522 seconds. In contrast, the proposed algorithm averaged 823 seconds per trip, and Baseline 2, an autonomous energy-efficient controller with no lane changes, resulted in the longest trip time of 1,047 seconds. These differences in longitudinal behavior are reflected in the velocity profiles shown in Fig. 7.
As a result, the proposed algorithm achieved a 39% reduction in motion energy consumption compared to Baseline 1, averaging $580$ Wh versus $808$ Wh. Furthermore, it achieved a 24% reduction in total energy consumption (including auxiliary power usage) compared to Baseline 1 on average. While one might argue that the energy savings are simply due to the vehicle driving more slowly, the proposed algorithm actively computed energy-efficient velocity profiles and lane selections based on an energy cost function. No lower bound was imposed on the velocity profile, allowing the algorithm to explore the energy-optimal behavior of the electric vehicle. Although Baseline 2 exhibited similar motion energy consumption to the proposed algorithm (within 2% on average), its significantly longer trip time led to higher overall energy usage due to increased auxiliary power consumption, resulting in only a 7% reduction on average. Ultimately, the proposed algorithm demonstrated the lowest total energy consumption as defined by the cost function. These findings raise compelling questions: Is slower driving always more energy-efficient for electric vehicles? And if not, what is the “sweet spot” for speed to balance energy efficiency and trip time? We leave these questions to be explored in future work.
Figure 6: Experiment Results: (a) Total motion energy consumption, calculated by subtracting auxiliary power usage from the overall energy consumption; (b) Overall energy consumption.
@@ VII Conclusion
This work presents energy-efficient lane change planning algorithms. By allowing flexible velocity profiles within legal speed limits, we explored the potential for minimizing energy consumption through intelligent lane selection and speed planning, enabled by I2V communication. Experimental results demonstrate that the proposed algorithm achieves significant reductions in motion energy consumption – up to 39% compared to a human driver. Notably, this was achieved without enforcing a travel time constraint, revealing the algorithm’s ability to discover energy-optimal behavior through real-time decision-making.
While the proposed method results in longer trip times than human driving, it highlights the trade-off between energy efficiency and travel time. Additionally, the comparison with a baseline controller that performs no lane changes underscores the importance of intelligent lane selection in reducing total energy use. These findings open up interesting avenues for future research, including the exploration of optimal speed “sweet spots” for electric vehicles, dynamic trade-off tuning between energy savings and travel time, and real-time adaptation to traffic conditions. The presented VIL platform offers a safe and effective tool for advancing the development and validation of sustainable autonomous driving technologies. While the proposed lane selection framework demonstrates promising energy savings on the tested urban route, several limitations should be acknowledged. The evaluation was conducted on a specific test route under moderate traffic conditions, and the results may not generalize directly to routes with significantly different topologies, elevation changes, or traffic densities. Future work will investigate the generalizability of the method across diverse environments and assess its robustness under varying traffic signal patterns and dynamic interactions with human drivers.
@@ VIII Acknowledgments
This research work presented herein is funded by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy under DE-AR0000791.

### v2

@@ VI-1 Experimental Setup
We evaluate the proposed approach using a Vehicle-in-the-Loop (VIL) system that integrates a Hyundai IONIQ 5 with CARLA simulation for realistic traffic and SPaT signals. Fig. 5 shows the virtual CARLA simulator map, the satellite image of the physical testing site, and the test vehicle (IONIQ 5). This setup enables robust evaluation of planning and control algorithms in a safe environment while using real energy measurements on the physical vehicle. See [11] for hardware details.
@@
Each experiment consists of three laps around the route (approximately 4 km). For comparison, we designed two baseline approaches and conducted experiments to collect the corresponding energy consumption measurements: (i) a human driver observing the virtual environment (Baseline 1), and (ii) an autonomous eco-driving controller that optimizes longitudinal speed without lane changes (Baseline 2), which is essentially a lane-keeping algorithm that optimizes vehicle behavior based on an energy stage cost.
@@ VI-2 Experiment Results
Fig. 6(a) presents the total motion energy consumption, calculated by subtracting auxiliary power usage from the overall energy consumption, while Fig. 6(b) shows the total energy usage including auxiliary loads. The human driver (Baseline 1) tended to drive closer to the speed limit and changed lanes as needed, resulting in the shortest average trip time of 522 seconds. In contrast, the proposed algorithm averaged 823 seconds per trip, and Baseline 2 resulted in the longest trip time of 1,047 seconds. These differences in longitudinal behavior are reflected in the velocity profiles shown in Fig. 7.
As a result, the proposed algorithm achieved a 39% reduction in motion energy consumption compared to Baseline 1, averaging $580$ Wh versus $808$ Wh. Furthermore, it achieved a 24% reduction in total energy consumption (including auxiliary power usage) compared to Baseline 1 on average. While one might argue that the energy savings are simply due to the vehicle driving more slowly, the proposed algorithm actively computed energy-efficient velocity profiles and lane selections based on an energy cost function. No lower bound was imposed on the velocity profile, allowing the algorithm to explore the energy-optimal behavior of the electric vehicle. Although Baseline 2 exhibited similar motion energy consumption to the proposed algorithm (within 2% on average), its significantly longer trip time led to higher overall energy usage due to increased auxiliary power consumption, resulting in only a 7% reduction on average. Ultimately, the proposed algorithm demonstrated the lowest total energy consumption as defined by the cost function.
Figure 6: Experiment Results: (a) Total motion energy consumption (Overall energy consumption - auxiliary power usage); (b) Overall energy consumption.
@@ VII Conclusion
This work presented an energy-aware lane selection framework for connected electric vehicles that integrates lane-change decisions with longitudinal speed optimization using V2I communication. Vehicle-in-the-loop experiments with a Hyundai IONIQ5 demonstrated up to 39% reduction in motion energy consumption compared to human driving and 7% savings relative to a lane-keeping eco-driving baseline, without imposing explicit travel-time constraints.
The results highlight the trade-off between energy efficiency and trip time, and show that lane selection is critical in reducing overall energy use for battery electric vehicles. While demonstrated on a specific route and traffic scenario, the framework illustrates how connectivity-enabled planning can improve the sustainability of autonomous urban driving.
Future work will investigate the generalizability of the method across diverse environments and assess its robustness under varying traffic signal patterns and dynamic interactions with human drivers.
@@ VIII Acknowledgments
This research work presented herein is funded by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy under DE-AR0000791. We also extend our sincere gratitude to the Hyundai Motor Company’s Vehicle Control Development Center and Autonomous Driving Development Center for their valuable technical assistance and support for the experimental vehicle.